# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v1

@@ IV Realizing negative inductance through quench
Our proposal thus requires a mechanism to locally address the Josephson junctions in a way to create connected regions with inverted inductance. We note that so-called tunable $0-\pi$ junctions are a subject of ongoing research and have multiple proposals for their implementation, including the theoretical proposal about Josephson junctions with a high-spin magnetic impurity sandwiched between two superconductors [89], experimental realization in ballistic Dirac semimetal Josephson junctions [90] and another experimental realization in Indium antimonide (InSb) two dimensional electron gases [91].
@@
For the matrix given in Eq. (23), we can explicitly compute a required current profile that needs to be applied in order to swap the phases by ±$\pm\pi$ within a connected part of the 1D chain. Blue dots in Fig. 4 show that if $\Delta l\ll\Delta x$, $\alpha$ is almost perfectly diagonal, such that there is a one-to-one correspondence between the desired target phase profile and the required current profile that has to be applied to the individual loops. Note however that the onset of non-local effects may play a role even for (moderately) small $\Delta l/\Delta x$ (orange and brown dots in Fig. 4), where we need to apply a slightly nonlocal current profile to have a sharp, local switch from positive to negative inductances within the 1D chain.
To conclude this section, we point out similarities and differences to previous proposals creating analog event horizons and Hawking radiation in solid state systems. Specifically, the analog Hawking radiation previously predicted in Josephson junction arrays [26, 27] relies on the creation of solitons, where within the finite extension of the soliton (over many lattice sites), the inductance can likewise be regarded as effectively negative within the soliton profile. The procedure we produce on the other hand involves creating ±$\pm\pi$ shifts on a connected series of junctions, which in essence corresponds to a tightly packed generation of “half”-solitons with the width of a single lattice site. Our system is thus arguably much less stable. But importantly, in conjunction with the nonreciprocity provided by the gyrators, it allows for the explicit realization of arbitrary (on-demand) spacetime geometries, and in addition, the explicit creation of wormholes with (in principle) distinguishable black and white hole horizons. As for the evolution of this highly unstable system beyond transient times, we provide some thoughts in Sec. VI. Another important difference is that we here have in principle perfect control on the position of the horizon, contrary to the solitons studied in Ref. [26, 27], which are autonomously moving with respect to the lab frame. Moreover, we can create an event horizon with the precision of a single lattice site. This is in contrast to all other analog Hawking radiation proposals (cold atoms, circuits, tilted Weyl semimetals), where the analog metric always changes over a finite width (a healing length). This is an interesting caveat specifically regarding the existence of a well-defined Hawking temperature as already discussed in Sec. II.
@@
The system we will study here has fifty sites ($J=50$) with the wormhole located between $j_{0}=15$ and $j_{1}=35$. For convenience, we impose periodic boundaries in position space (i.e., the 1D chain is closed into a loop). The parameters of the system are $|E^{\prime}_{L}|=|E_{L}|=3.42E_{C}$, $G=2.4$ and $M=3*10^{-4}E_{C}$. As already foreshadowed, in each of the following figures of correlation functions the result could be obtained by both diagonalization and the generalized Klich determinant. The time steps in the following plots are $\delta t=0.031E_{C}^{-1}$ for the circuit with nearest neighbor coupling and $\delta t=0.31E_{C}^{-1}$ for the next-to-nearest neighbor coupling.
@@
First let us examine the plots of correlation functions for the circuit with only nearest neighbor coupling, Fig. 5. We see that both phase and charge quantum fluctuations diverge after the quench. The system has two boundaries (red dotted lines) between the wormhole and the normal regions, that act as the horizons (located at $j_{0}$ and $j_{1}$). If we initialize low energy excitations (for this circuit this means excitations near $k=0$) in the region with overtilted dispersion, they move from the left horizon to the right horizon, hence we label them as the black and the white hole horizon respectively. The black hole horizon radiates away with time, as can be inferred by the accumulation of quantum fluctuations near it with time (which indicates presence of radiation), while the white hole horizon does not radiate. This explicitly confirms the previously discussed fact that for the nearest neighbour system, black and white hole horizons are distinguishable. The wormhole region also starts radiating due to the presence of the exceptional point (as likewise already discussed in Sec. III). We observer that the decay in the wormhole interior is much faster than the one induced by the presence of the horizons, which leads to the correlation functions growing much more rapidly inside the wormhole, before the black hole horizon show any appreciable decay. Overall, the above thus illustrates the possibility to create lattice simulations of wormholes with distinguishable black and white hole horizons, at the cost of a strong instability of the wormhole interior.
Now, let us focus on the circuit where Josephson junctions connect next nearest neighbors. Low-energy excitations in this circuit can be either near $k=0$ or $k=\pm\pi$. For excitations (inside the wormhole region) near $k=0$, similar to previous circuit the left boundary ($j_{0}$) acts as the black hole horizon and the right one ($j_{1}$) as the white hole. In contrast, for excitations near $k=\pm\pi$ the left boundary acts as the white hole horizon and right boundary as the black hole. This symmetry is also reflected in the correlation plots in Fig. 6, where we can observe both horizons radiating away identically. Also the only instability in this circuit is due to the horizons, leading to a much slower collapse.
@@ VI Perspective on wormhole evaporation over long times
To conclude this section, we consider the very hypothetical case of negligble coupling to the environment, as it reveals an interesting additional idea which we pursue in the remainder of this work. If we assume that the build-up of quantum fluctuations of the phase could progress unimpeded over longer times, the nonlinearity of the Josephson energies results not in classical fluctuations of the effective inductance (as in the previous paragraph), but in a system that has to be interpreted as being in a quantum superposition of different effective inductances – a form of quantum superposition of the spacetime geometry itself. However, within the above considered setting, this effect is likely not to survive for very long due to environment-induced decoherence of the quantum fluctuations, and even for a highly protected system, it would be hard to analyse the effect in an unambiguous way. In what follows now, we provide a modification of this idea, not relying on instabilities (and thus not requiring any ultra-fast quenches), but instead on more general notions of Josephson junctions with multiple minimas. This endeavour will on the one hand introduce the notion of a quantum inductance, and with it, a likely more stable version of a quantum superposition of the analog spacetime.
@@ VII Bistable Josephson effect and quantum spacetime geometry
As announced, in this section we propose a possible way to create and observe superposition of two different spacetime geometries using quantum circuits. As already hinted at in the introduction, the idea that spacetime geometries can exist as a superposition within actual gravity is exciting, but at the same time highly contentious [46].
Note moreover, that the currently most debated possibility is the observation of entanglement created between two quantum systems via the gravitational force, hence proving quantum nature of gravitation [42, 43]. Since gravity, in part, is a theory of structure of spacetime, it being quantum in nature may imply that spacetime can exist in quantum superposition. The notion that we introduce now is insofar related, as it allows for the formulation of a quantum superposition of analog spacetime geometry, by exploiting internal degrees of freedom of general non-linear inductor elements.
The central idea can be formulated as follows. Previously, the local signal speed in a circuit composed of Josephson junctions, is uniquely defined by the curvature of the minimum of the Josephson potential. However, if we use a potential with more than one minimum, where – crucially – each minimum has a different curvature, then they will have different signal speed. Since these two minimas can be considered a quantum double well, it is possible to prepare a state where one occupies a superposition of the two minima, ultimately allowing us to create a quantum superposition of signals travelling simultaneously at multiple different speeds.
@@
with $|\delta|<1/2$. This function has two minimas, whose curvatures are $\sim 4(1\pm 2\delta)$. We note that while the parallel shunt of the above three different junction types is very instructive to demonstrate the principle, the actual experimental fabrication of a potential of the form of Eq. (27) may be simplified. For instance, it is expected [115] that the setup realizing the term $\sim\cos\left(3\phi\right)$ [Fig. 8(b)] has enough free parameters to implement the potential without the additional inclusion of a $\cos(\phi)$ and a $\cos(\phi)$ junction.
@@
At any rate, the above potential has the sought-after minima with different curvatures. This circuit element is therefore at low-energies described as a type of quantum inductor, which can exhibit superpositions of inductance values. Using such an element to connect nodes $j$ and $j-1$, this element can be described by the energy
@@
where the relative difference of inductances is given by $\delta E_{L}/E_{L}=2\delta$ [with $\delta$ as it appears in Eq. (27)], and the pseudo-spin $\sigma^{z}$ represents the two distinct quantum well states of the inductor, $\left|0_{j}\right>$ and $\left|1_{j}\right>$, having eigenvalue $-1$ and $+1$, respectively. We note that this description is of course a crude approximation. In particular, finite capacitances will likely induce transitions $\sim\sigma^{x}_{j}$. While we for now discard this possibility for simplicity, we will outline in the outlook, that the inclusion of coherent spin-flip processes (or even nonlocal spin-spin interactions) by no means pose a danger to the proposal outlined here, but to the contrary, render it even more interesting (as they may give rise to a rudimentary simulation of an analog quantum spacetime with an intrinsic coherent dynamics of its own).
The full proposal to realize and measure superposition of spacetime geometries now works as follows. Consider the circuit in Fig. 9 consisting of three regions, two regions where the superconducting nodes are connected by Josephson junctions and a third one sandwiched between the the other two where superconducting nodes are connected by the newly introduced quantum inductors. This system is represented by the following Hamiltonian
@@
To measure this entanglement we will exploit the fact that the two states of the asymmetric potential correspond to different signal speeds, present in the sandwiched region. Let us now initialize a wave packet in the trivial region to the left. Assuming for concreteness a signal with large amplitude in $\phi$, we can simply describe it with a wave function $\phi(x-ut)$ (using a continuous space representation $x=\Delta xj$, valid if the wave packet is large with respect to the lattice size $\Delta x$), centered around a position on the left, and moving towards the region with the quantum inductors. The propagation speed here is $u=\Delta x\sqrt{E_{C}E_{L}^{\prime}}$ Within this region, the signal will now propagate with two different speeds. If the quantum inductor region is large enough with respect to the speed difference and the wave packet size, we will create a fully entangled state. The transition from the incoming wave packed to the left of the quantum inductor region, and the outgoing wave packet to the right of the quantum inductor region can be described as
@@
The positions $x_{0}$ and $x_{1}>x_{0}$ (i.e., the wave packet $\phi(x+x_{1}-ut)$ is the further advanced one) express the different positions of the outgoing wave packets due to the different times of flight within the quantum inductor region. If the overlap vanishes, $\phi(x+x_{0}-ut)\phi(x+x_{1}-ut)\approx 0$, the entanglement is maximal.
@@
where in the last line, the updated shift $\widetilde{x}_{1}$ takes into account the distance travelled up to and after collision with the hard wall. As can be seen on the last line of above equation, the further advanced package now travels backwards ($+ut$ instead of $-ut$), and is thus bound to interfere with the delayed wave packet (which still travels forward, $-ut$). In this setup, interference is constructive for the $\left|\tilde{0}\right>$ branch, and destructive for the $\left|\tilde{1}\right>$ branch. Hence, by projective measurement of the signal either directly after passing the quantum inductor region, or measuring after the succession of Hadamard gate and reflection, the quantum nature of the spacetime geometry can be unequivocally probed.
@@ VIII Conclusions and outlook
In this paper we have proposed using quantum circuits to implement two aspects of quantum gravity: 1) analog horizons and 2) quantum superposition of spacetime geometries. Even though the former idea is a well explored concept in solid state systems in general, we here explicitly demonstrate the capacity of superconducting circuits to emulate arbitrary spacetime configurations. By identifying a minimal fundamental set of necessary circuit elements, we further unravel a number of surprising findings pertinent to lattice systems, which allow in particular to create horizons with trans-Planckian curvature, allowing us to disentangle the effect of curvature of background metric from the effect of a horizon, which as far as we know has not been done in any other proposal of analog horizons. To implement the analog horizons in a quantum circuit, we proposed a general way of creating negative inductances with Josephson junctions transiently driven by nearby current loops. We further uncovered two subtly distinct ways to engineer a region with overtilted dispersion relation, either with inductive coupling of nearest and next-to-nearest neighbour nodes. While for the latter, horizons have a combination black and white hole character (as is the most usual case in lattices), the former [Eq.(III)] hosts two distinct kinds of horizons, one that acts purely as a black hole horizon and the other as a white hole horizon. For this second type of circuit, it is not only the horizons, but also the region between the horizons (wormhole interior) that are unstable due to presence of an exceptional point. This instability also contributes to the observed radiation (in addition to the evaporation of the horizons) as can be confirmed by tracking the quantum fluctuations of phase difference and conjugate charge with time (Fig. 5). In addition to Bogoliubov diagonalization, we used an extension of Klich’s trace-determinant formula to obtain the correlation functions numerically. The proof of this extension remains elusive for a general operator, but in the parameter regime where the generalised Bogoliubov transformation is applicable, the results of both the numerical methods agree.
@@
Our second proposal concerns the controversial idea of superposition of spacetime geometries. This idea was shown to emerge as a nontrivial extension to the above negative inductance realization. The emulation of such geometries relies on engineering junction elements with an energy-phase relationship that exhibit an asymmetric double well potential (Eq.(27)), so-called multi-stable Josephson junctions. The different curvatures of the two minima of the potential give us the two inductances that act as the basis states for the quantum inductance. We proposed a setup where a classical signal traverses the quantum spacetime region, resulting in a superposition of two different signal speeds. The superposition can be measured by making correlated measurements of the quantum inductance and the signal moving through the circuit, or by means of a delayed choice experiment.
@@
In this paper we have either used circuit elements that are still being developed experimentally and therefore are not standard in cQED (such as quantum gyrators), or introduced new effective circuit elements (such as regular or multistable Josephson junctions), which, even though experimentally tested in existing works, are here proposed to be operated in an exotic regime. We here therefore point at feasible small scale experimental tests that can act as proofs of principle for the basic physical principles behind the considered physics. For instance, our proposal of using current loops to invert the effective inductance of a Josephson junction can be tested on a single transmon, where the qubit frequency becomes imaginary $\sqrt{E_{L}E_{C}}\to\sqrt{-E_{L}E_{C}}$ for transient times. The resulting accumulation phase or charge quantum fluctuations (or limits thereof due to coupling to the environment) can thus be tested in an “imaginary” qubit device. Such a study might be also of some fundamental interest, when extending the description of the junction beyond the quadratic approximation, where the expansion around the energy minimum or the energy maximum can be regarded as real and imaginary twins of nonlinear dynamics. For the case of quantum inductors, one can likewise consider a simple circuit with a single quantum inductor and a capacitor, where the low energy excitations induced in the two minima will have different oscillation frequency. Finally, we comment on extensions towards a more fully-fletched simulation of analog quantum spacetimes. Namely, in our treatment in the main text, the states of the quantum inductors had as of now no intrinsic dynamics. While this simplified treatment is the simplest from a theoretical stand point, it requires a tuning to very extreme ratios of small charging energy compared to the junction energy scale. For more moderate ratios, it cannot be excluded that the basis states of the quantum inductors receive an intrinsic dynamics, for example a tunnel coupling between the two minimas, $\sigma_{j}^{x}$, or non-local couplings such as $\sigma_{j}^{\alpha}\sigma_{j^{\prime}}^{\beta}$ ($\alpha,\beta=x,y,z$), which have to be added in the Hamiltonian in Eq.(29).

### v2

@@ IV Realizing negative inductance through quench
Our proposal, thus, requires a mechanism to locally address the Josephson junctions in a way to create connected regions with inverted inductance. We note that so-called tunable ${0-\pi}$ junctions are a subject of ongoing research and have multiple proposals for their implementation, including the theoretical proposal about Josephson junctions with a high-spin magnetic impurity sandwiched between two superconductors [97], experimental realization in ballistic Dirac semimetal Josephson junctions [98] and another experimental realization in Indium antimonide (InSb) two dimensional electron gases [99].
@@
For the matrix given in Eq. (23), we can explicitly compute a required current profile that needs to be applied in order to swap the phases by ±$\pm\pi$ within a connected part of the 1D chain. Blue dots in Fig. 4 show that if $\Delta l\ll\Delta x$, $\alpha$ is almost perfectly diagonal, such that there is a one-to-one correspondence between the desired target phase profile and the required current profile that has to be applied to the individual loops. Note however that the onset of non-local effects may play a role even for (moderately) small $\Delta l/\Delta x$ (orange and brown dots in Fig. 4), where we need to apply a slightly nonlocal current profile to have a sharp, local switch from positive to negative inductances within the 1D chain. To conclude this section, we point out similarities and differences to previous proposals creating analog event horizons and Hawking radiation in solid state systems. Specifically, the analog Hawking radiation previously predicted in Josephson junction arrays [32, 33] relies on the creation of solitons, where within the finite extension of the soliton (over many lattice sites), the inductance can likewise be regarded as effectively negative within the soliton profile. The procedure we produce on the other hand involves creating ±$\pm\pi$ shifts on a connected series of junctions, which in essence corresponds to a tightly packed generation of “half”-solitons with the width of a single lattice site. Our system thus seems arguably much less stable. But first, as indicated already, the nonreciprocity provided by the gyrators allows for a stabilization of the system even with inverted inductors, due to an extra (positive) inductive contribution. Second, the extra tilting mechanism provided by the gyrators allows for the explicit realization of arbitrary (on-demand) spacetime geometries, and in addition, the explicit creation of wormholes with (in principle) distinguishable black and white hole horizons. As for the evolution of the unstable system beyond transient times, we provide some thoughts in Sec. VI. Another important difference is that we here have in principle perfect control on the position of the horizon, contrary to the solitons studied in Ref. [32, 33], which are autonomously moving with respect to the lab frame. Moreover, we can create an event horizon with the precision of a single lattice site. This is in contrast to all other analog Hawking radiation proposals (cold atoms, circuits, tilted Weyl semimetals), where the analog metric always changes over a finite width (a healing length). This is an interesting caveat specifically regarding the existence of a well-defined Hawking temperature, or more generally, how to make sense of horizons with diverging curvature, as already indicated in Sec. II.
@@
The system we will study here has fifty sites ($J=50$) with the wormhole located between $j_{0}=15$ and $j_{1}=35$. For convenience, we impose periodic boundaries in position space (i.e., the 1D chain is closed into a loop). The parameters of the system are $|E^{\prime}_{L}|=|E_{L}|=3.42E_{C}$, $G=2.4$ and $M=3*10^{-4}E_{C}$. As already pointed out, in each of the following figures of correlation functions the result could be obtained by both diagonalization and the generalized Klich determinant. The time steps in the following plots are $\delta t=0.031E_{C}^{-1}$ for the circuit with nearest neighbor coupling and $\delta t=0.31E_{C}^{-1}$ for the next-to-nearest neighbor coupling.
@@
First, let us examine the plots of correlation functions for the circuit with only nearest neighbor coupling, Fig. 5. We see that both phase and charge quantum fluctuations diverge after the quench. The system has two boundaries (red dotted lines) between the wormhole and the normal regions, that act as the horizons (located at $j_{0}$ and $j_{1}$). If we initialize low energy excitations (for this circuit this means excitations near $k=0$) in the region with overtilted dispersion, they move from the left horizon to the right horizon, hence we label them as the black and the white hole horizon respectively. The black hole horizon radiates away with time, as can be inferred by the accumulation of quantum fluctuations near it with time (which indicates presence of radiation), while the white hole horizon does not radiate. This explicitly confirms the previously discussed fact that for the nearest neighbour system, black and white hole horizons are distinguishable. The wormhole region also starts radiating due to the presence of the exceptional point (as likewise already discussed in Sec. III). We observe that the decay in the wormhole interior is much faster than the one induced by the presence of the horizons, which leads to the correlation functions growing much more rapidly inside the wormhole, before the black hole horizon show any appreciable decay. Overall, the above thus illustrates the possibility to create lattice simulations of wormholes with distinguishable black and white hole horizons, at the cost of a strong instability of the wormhole interior.
Now, let us focus on the circuit where Josephson junctions connect next nearest neighbors. Low-energy excitations in this circuit can be either near $k=0$ or $k=\pm\pi$. For excitations (inside the wormhole region) near $k=0$, similar to the previous circuit, the left boundary ($j_{0}$) acts as the black hole horizon and the right one ($j_{1}$) as the white hole. In contrast, for excitations near $k=\pm\pi$ the left boundary acts as the white hole horizon and right boundary as the black hole. This symmetry is also reflected in the correlation plots in Fig. 6, where we can observe both horizons radiating away identically. Also, the only instability in this circuit is due to the horizons, leading to a much slower collapse.
@@ VI Perspective on wormhole evaporation over long times
To conclude this section, we consider the very hypothetical case of negligible coupling to the environment, as it reveals an interesting additional idea which we pursue in the remainder of this work. If we assume that the build-up of quantum fluctuations of the phase could progress unimpeded over longer times, the nonlinearity of the Josephson energies results not in classical fluctuations of the effective inductance (as in the previous paragraph), but in a system that has to be interpreted as being in a quantum superposition of different effective inductances – a form of quantum superposition of the spacetime geometry itself. However, within the above considered setting, this effect is likely not to survive for very long due to environment-induced decoherence of the quantum fluctuations, and even for a highly protected system, it would be hard to analyse the effect in an unambiguous way. In what follows now, we provide a modification of this idea, not relying on instabilities (and thus not requiring any ultra-fast quenches), but instead on more general notions of Josephson junctions with multiple minimas. This endeavour will on the one hand introduce the notion of a quantum inductance, and with it, a likely more stable version of a quantum superposition of the analog spacetime.
@@ VII Bistable Josephson effect and quantum spacetime geometry
As announced, in this section we propose a possible way to create and observe superpositions of two different spacetime geometries using quantum circuits. As already hinted at in the introduction, the idea that spacetime geometries can exist as a superposition within actual gravity is exciting, but at the same time highly contentious [54].
Note, moreover, that the currently most debated possibility is the observation of entanglement created between two quantum systems via the gravitational force, hence proving the quantum nature of gravitation [50, 51]. Since gravity, in part, is a theory of the structure of spacetime, it being quantum in nature may imply that spacetime can exist in quantum superposition. The notion that we introduce now is insofar related, as it allows for the formulation of a quantum superposition of analog spacetime geometry, by exploiting internal degrees of freedom of general non-linear inductor elements.
The central idea can be formulated as follows. Previously, the local signal speed in a circuit composed of Josephson junctions, is uniquely defined by the curvature of the minimum of the Josephson potential. However, if we use a potential with more than one minimum, where – crucially – each minimum has a different curvature, they will have different signal speed. Since these two minima can be considered a quantum double well, it is possible to prepare a state where one occupies a superposition of the two minima, ultimately allowing us to create a quantum superposition of signals travelling simultaneously at multiple different speeds.
@@
with $|\delta|<1/2$. This function has two minima, whose curvatures are $\sim 4(1\pm 2\delta)$. We note that while the parallel shunt of the above three different junction types is very instructive to demonstrate the principle, the actual experimental fabrication of a potential of the form of Eq. (27) may be simplified. For instance, it is expected [123] that the setup realizing the term $\sim\cos\left(3\phi\right)$ [Fig. 8(b)] has enough free parameters to implement the potential without the additional inclusion of a $\cos(\phi)$ and a $\cos(2\phi)$ junction.
@@
At any rate, the above potential has the sought-after minima with different curvatures. This circuit element is therefore at low energies described as a type of quantum inductor, which can exhibit superpositions of different inductance values. Using such an element to connect nodes $j$ and $j-1$, this element can be described by the energy
@@
where the relative difference of the inductances is given by $\delta E_{L}/E_{L}=2\delta$ [with $\delta$ as it appears in Eq. (27)], and the pseudo-spin $\sigma^{z}$ represents the two distinct quantum well states of the inductor, $\left|0_{j}\right>$ and $\left|1_{j}\right>$, having eigenvalue $-1$ and $+1$, respectively. We note that this description is of course a crude approximation. In particular, finite capacitances will likely induce transitions $\sim\sigma^{x}_{j}$. While we for now discard this possibility for simplicity, we will outline in the outlook, that the inclusion of coherent spin-flip processes (or even nonlocal spin-spin interactions) by no means pose a danger to the proposal outlined here, but to the contrary, render it even more interesting (as they may give rise to a rudimentary simulation of an analog quantum spacetime with an intrinsic coherent dynamics of its own).
The full proposal to realize and measure superpositions of spacetime geometries now works as follows. Consider the circuit in Fig. 9 consisting of three regions, two regions where the superconducting nodes are connected by Josephson junctions and a third one sandwiched between the the other two where superconducting nodes are connected by the newly introduced quantum inductors. This system is represented by the following Hamiltonian
@@
To measure this entanglement we will exploit the fact that the two states of the asymmetric potential correspond to different signal speeds, present in the sandwiched region. Let us now initialize a wave packet in the trivial region to the left. Assuming for concreteness a signal with large amplitude in $\phi$, we can simply describe it with a wave function $\phi(x-ut)$ (using a continuous space representation $x=\Delta xj$, valid if the wave packet is large with respect to the lattice size $\Delta x$), centered around a position on the left, and moving towards the region with the quantum inductors. The propagation speed here is $u=\Delta x\sqrt{E_{C}E_{L}^{\prime}}$. Within this region, the signal will now propagate with two different speeds. If the quantum inductor region is large enough with respect to the speed difference and the wave packet size, we will create a fully entangled state. The transition from the incoming wave packed to the left of the quantum inductor region, and the outgoing wave packet to the right of the quantum inductor region can be described as
@@
The positions $x_{0}$ and $x_{1}>x_{0}$ (i.e., the wave packet ${\phi(x+x_{1}-ut)}$ is the further advanced one) express the different positions of the outgoing wave packets due to the different times of flight within the quantum inductor region. If the overlap vanishes, i.e. if ${\phi(x+x_{0}-ut)\phi(x+x_{1}-ut)\approx 0}$, the entanglement is maximal.
@@
where in the last line, the updated shift $\widetilde{x}_{1}$ takes into account the distance travelled up to and after collision with the hard wall. As can be seen on the last line of the above equation, the further advanced package now travels backwards ($+ut$ instead of $-ut$), and is thus bound to interfere with the delayed wave packet (which still travels forward, $-ut$). In this setup, interference is constructive for the $\left|\tilde{0}\right>$ branch, and destructive for the $\left|\tilde{1}\right>$ branch. Hence, by projective measurement of the signal either directly after passing the quantum inductor region, or measuring after the succession of Hadamard gate and reflection, the quantum nature of the spacetime geometry can be unequivocally probed.
@@ VIII Conclusions and outlook
In this paper we have proposed using quantum circuits to implement two aspects of quantum gravity: 1) analog horizons and 2) quantum superpositions of spacetime geometries. Even though the former idea is a well explored concept in solid state systems in general, we here explicitly demonstrate the capacity of superconducting circuits to emulate arbitrary spacetime configurations. By identifying a minimal fundamental set of necessary circuit elements, we further unravel a number of surprising findings pertinent to lattice systems, which allow in particular to create horizons with trans-Planckian curvature, allowing us to disentangle the effect of curvature of background metric from the effect of a horizon – and in doing so, exploring the extreme regime of horizons with diverging curvatures, which as far as we know has not been done in any existing proposal. To implement the analog horizons in a quantum circuit, we proposed a general way of creating negative inductances with Josephson junctions transiently driven by nearby current loops. We further uncovered two subtly distinct ways to engineer a region with overtilted dispersion relation, either with inductive coupling of nearest or next-to-nearest neighbour nodes. While for the latter, horizons have a combination of black and white hole character (as is the most usual case in lattices), the former [Eq.(III)] hosts two distinct kinds of horizons, one that acts purely as a black hole horizon and the other as a white hole horizon. For this second type of circuit, it is not only the horizons, but also the region between the horizons (wormhole interior) that is unstable due to the presence of an exceptional point. This instability also contributes to the observed radiation (in addition to the evaporation of the horizons) as can be confirmed by tracking the quantum fluctuations of phase difference and conjugate charge with time (Fig. 5). In addition to Bogoliubov diagonalization, we used an extension of Klich’s trace-determinant formula to obtain the correlation functions numerically. The proof of this extension remains elusive for a general operator, but in the parameter regime where the generalised Bogoliubov transformation is applicable, the results of both numerical methods agree.
@@
Our second proposal concerns the controversial idea of superpositions of spacetime geometries. This idea was shown to emerge as a nontrivial extension to the above negative inductance realization. The emulation of such geometries relies on engineering junction elements with an energy-phase relationship that exhibits an asymmetric double well potential (Eq.(27)), so-called multi-stable Josephson junctions. The different curvatures of the two minima of the potential give us the two inductances that act as the basis states for the quantum inductance. We proposed a setup where a classical signal traverses the quantum spacetime region, resulting in a superposition of two different signal speeds. The superposition can be measured by making correlated measurements of the quantum inductance and the signal moving through the circuit, or by means of a delayed choice experiment.
@@
In this paper, in addition to standard circuit elements, we have also included elements that are still being developed experimentally (such as quantum gyrators), or introduced new effective circuit elements (based on regular or multistable Josephson junctions), which, even though experimentally tested in existing works, are here proposed to be operated in previously unexplored regimes. We here therefore point at feasible small scale experimental tests that can act as proofs of principle for the basic physical principles behind the considered physics. For instance, our proposal of using current loops to invert the effective inductance of a Josephson junction can be tested on a single transmon, where the qubit frequency becomes imaginary $\sqrt{E_{J}E_{C}}\to\sqrt{-E_{J}E_{C}}$ for transient times. The resulting accumulation of phase or charge quantum fluctuations (or limits thereof due to coupling to the environment) can thus be tested in an “imaginary” qubit device. Such a study might be also of some fundamental interest, when extending the description of the junction beyond the quadratic approximation, where the expansion around the energy minimum or the energy maximum can be regarded as real and imaginary twins of nonlinear dynamics – ready to be explored in future works. For the case of quantum inductors, one can likewise consider a simple circuit with either one or a few quantum inductors in series, where the low energy excitations induced in the two minima will have different oscillation frequency – allowing for the engineering of (nearly) harmonic oscillators in a quantum superposition of two resonance frequencies.
Finally, we comment on extensions towards a more fully-fletched simulation of analog quantum spacetimes. Namely, in our treatment in the main text, the states of the quantum inductors had, as of now, no intrinsic dynamics. While this is the simplest possilbe treatment from a theoretical stand point, in experiment it would require a tuning to very extreme ratios of small charging energy compared to the junction energy scale. For more moderate ratios, it cannot be excluded that the basis states of the quantum inductors receive an intrinsic dynamics, for example a tunnel coupling between the two minima, $\sigma_{j}^{x}$, or non-local couplings such as $\sigma_{j}^{\alpha}\sigma_{j^{\prime}}^{\beta}$ ($\alpha,\beta=x,y,z$), which have to be added in the Hamiltonian in Eq.(29). In this regard, note that this specific quantum spacetime proposal is in line with currently existing plans to engineer quantum computing hardware by means of flux qubits based on multi-stable Josephson junctions [100, 101, 102]. Crucially, however, while the above described type of crosstalk is obviously detrimental for the performance of the hardware when used for standard quantum information applications, for the here considered analog quantum gravity simulation, such imperfections are by no means a disadvantage – to the contrary, they may open up a new road to explore analog quantum gravity where the spacetime metric has its quantum coherent dynamics.